Kayla uses her credit card to purchase a new television for $487.89. She can pay off up to $175 per month. The card has an annual rate of 23.5% compounded monthly. How much will she pay in interest?
Answer by Theo(13342)

you can use the following online calculator to answer your question.

http://www.arachnoid.com/lutusp/finance.html

you would make the folowing entries:

Present Value = 487.89
Future Value = 0
Number of Payments is left blank because you want to solve for that.
Payment Amount = -175.00
Interest Rate per period = 23.5% per year divided by 12 = 1.958333% per month.
Payment At is at End.

click on the NP button and it will tell you the number of months required for the loan to be fully paid off.

the calculator will show you that it will take 2.89 months for you to fully pay off the loan of $487.99 when you are paying it off at $175.00 at the end of each momth and the interest rate if 1.958333% per month.

now that you know how long it takes for you to pay it off, you can determine what the interest amount is that you paid.

multiply your payment of 175 per month by 2.89 months to get a total payment of $505.75.

your loan was for $487.89.

the difference is the amount of interest you paid.

that will be $505.75 - $487.89 = $17.86.\

you paid $17.86 in interest.

17.86 / 487.89 is equal to roughly 3.7% of the purchase price.

that's the total interest you paid on the loan.

why so low?

because you paid the loan off in less than 3 months.

if you had taken 2 years to pay off the lona, you would have had to pay more interest.

i'll do the calculations for you so you can see.

Present Value is 487.89
Future Value is 0
Number of Payments is 24
Payment Amount is what you want to find.
Interest Rate per period = 1.958333%
Payment At End of month

click on PMT and you will need to be paying 25.67 per month in order to completely pay off the loan in 24 months.

how much interest would you have paid?

24 * 25.67 = 616.08

616.08 - 487.89 = 128.19

128.19 / 487.89 = 26%.

if you had paid the loan off in 24 months, you would have paid 128.19 in interest which was equivalent to 26% of the initial purchase price.

the fact that you paid the loan off so quickly is why you were not charged a higher percentage of the initial purchase price.